# Unifying Specialized Visual Encoders for Video Language Models

### v1

@@ Abstract
The recent advent of Large Language Models (LLMs) has ushered sophisticated reasoning capabilities into the realm of video through Video Large Language Models (VideoLLMs). However, VideoLLMs currently rely on a single vision encoder for all of their visual processing, which limits the amount and type of visual information that can be conveyed to the LLM. Our method, MERV, Multi-Encoder Representation of Videos, instead leverages multiple frozen visual encoders to create a unified representation of a video, providing the VideoLLM with a comprehensive set of specialized visual knowledge. Spatio-temporally aligning the features from each encoder allows us to tackle a wider range of open-ended and multiple-choice video understanding questions and outperform prior state-of-the-art works. MERV is up to 3.7% better in accuracy than Video-LLaVA across the standard suite video understanding benchmarks, while also having a better Video-ChatGPT score. We also improve upon SeViLA, the previous best on zero-shot Perception Test accuracy, by 2.2%. MERV introduces minimal extra parameters and trains faster than equivalent single-encoder methods while parallelizing the visual processing. Finally, we provide qualitative evidence that MERV successfully captures domain knowledge from each of its encoders. Our results offer promising directions in utilizing multiple vision encoders for comprehensive video understanding.
@@ 1 Introduction
Inspired by the sophisticated reasoning abilities of recent Large Language Models (LLMs) (Chiang et al., 2023; Chowdhery et al., 2023; OpenAI, 2023), researchers have focused on using them in many other domains to great success. The video counterparts, known as Video Large Language Models (VideoLLMs) (Bain et al., 2021; Li et al., 2023c; Lin et al., 2023a; Luo et al., 2023; Maaz et al., 2023; Yu et al., 2024), connect pretrained vision encoders to LLMs by training a modality bridge from the vision space to the language space, allowing for reasoning to happen in the highly expressive language domain.
Most multimodal LLMs, such as LLaVA (Liu et al., 2023) for images and Video-LLaVA (Lin et al., 2023a) for videos, opt for contrastively pretrained encoders like CLIP (Radford et al., 2021) and LanguageBind (Zhu et al., 2023). Their vision-language pretraining naturally lends itself as a bridge between the vision input and the LLM, circumventing the need to train heavy vision-language alignment modules like a QFormer (Li et al., 2022). These encoders are almost always pretrained separately and vary in architecture, training data, and optimization strategy. Consequently, the features extracted by these encoders exhibit unique characteristics, each with inherent strengths and limitations. Contrastive encoders like CLIP (Radford et al., 2021) may be better suited with their multimodal semantic alignment, but are inferior to models such as DINOv2 (Oquab et al., 2023) at fine-grained object level understanding. They also fail to take advantage of models trained specifically on videos, such as ViViT (Arnab et al., 2021). Despite this clear tension between vision backbones, previous research in VideoLLMs has relied on only one vision encoder for visual processing as one was thought to be sufficient for visual understanding, and already difficult enough to achieve vision-language alignment with. Any more encoders was unnecessary and not an effective tradeoff of runtime for compute.
@@
Figure 1: Different visual experts exhibit individual strengths. We show some examples where one single encoder model is the only model to correctly answer Perception Test questions (Pătrăucean et al., 2023).
@@
In this paper, we argue that this choice to not use multiple encoders in existing VideoLLMs unnecessarily restricts their capabilities. For example, in Figure 1 we can see cases where only one of four different single-encoder models answers a given question correctly. While simple scene descriptions can be answered by image-level models, other questions require temporal and action-level comprehension, benefiting from features encoded with video models like ViViT (Arnab et al., 2021). Consequently, the reasoning capabilities of these VideoLLMs are directly limited by the inherent weaknesses of their respective pretrained encoders. Therefore, employing multiple encoders could allow us to complement one encoder’s weaknesses with another encoder’s strengths. The wide adoption of the LLaVA paradigm is also indication that vision-language alignment is simple to achieve, even without language-aware vision models.
@@
We propose MERV, a Multi-Encoder Representation of Videos, as a new method for integrating multiple visual encoders into a single VideoLLM using a cross-attentive encoder mixer for fusing representations. We introduce a spatio-temporally aligned representation for mixing the information from multiple types of visual encoders. Given the computational complexity of video tasks, we carefully experiment with optimization strategies and parallelizing the visual experts, allowing us to combine four distinct visual encoders with minimal computational overhead. Our frozen method outperforms all individual encoder methods, is up to 3.7% better than prior works (Lin et al., 2023a) on video reasoning benchmarks, i.e., from 47.1% to 50.8% on ActivityNet-QA (Yu et al., 2019), and on par with the state-of-the-art (Yu et al., 2024) on Perception Test (Pătrăucean et al., 2023), a challenging perception and reasoning diagnostic for video models. Finetuning the full model improves MERV past SeViLA (Yu et al., 2024) by 2.2%, from 46.2% to 48.4%. Finally, we do a detailed qualitative study of our model’s capabilities on the Something-Something v2 dataset (Goyal et al., 2017). We show that MERV can accurately capture both contrastive encoders’ (Zhai et al., 2023; Zhu et al., 2023) strengths on general vision-language understanding, as well as ViViT’s (Arnab et al., 2021) specialty on temporally-sensitive tasks (e.g. distinguishing pushing left vs. right), without trading off performance between these specializations as single encoder models do. Our code and pretrained weights are available at https://github.com/princetonvisualai/merv.
@@ 2 Related Works
VideoLLMs build upon the powerful reasoning capabilities of LLMs by utilizing them as language decoders to enable instruction-followed video understanding. Key advancements include VideoChat (Li et al., 2023c) and Video-LLaMA (Zhang et al., 2023a) for chat-based video understanding, LLaMA-Adapter (Zhang et al., 2023b) for pre-alignment, Valley (Luo et al., 2023) with multilingual LLMs, InternVideo (Wang et al., 2022) with a dedicated video encoder training phase, and Video-ChatGPT (Maaz et al., 2023) combining video-adapted encoders with LLMs. GPT4Video (Wang et al., 2023b) supports video understanding and generation, while MovieChat (Song et al., 2023) focuses on long video comprehension. Models like Chat-UniVi (Jin et al., 2023) and LLaMA-VID (Li et al., 2023e) optimize token usage for video representation. Other notable models include Vamos (Wang et al., 2023a), which flexibly uses visual embeddings, action labels, and video captions as input; VideoChat2 (Li et al., 2023d), developed through three-stage progressive training; Video-LLaVA (Lin et al., 2023a), which aligns image and video representations before projecting them to the LLM space; and VideoPrism (Zhao et al., 2024), which also further trains a video encoder through masked distillation. Specialized models like VTimeLLM (Huang et al., 2023) focus on fine-grained video moment understanding and time-bound reasoning, while models like Elysium (Wang et al., 2024) and Merlin (Yu et al., 2023) can predict object trajectories. SeViLA (Yu et al., 2024) uses LLM for frame localizer of the video for multiple-choice tasks. Finally, recently LLaVA-Hound-DPO (Zhang et al., 2024) explored using DPO and a higher quality training set for better instruction following. Distinct from these aforementioned works, our approach centers on utilizing a diverse array of visual and especially video encoders, each with its own unique strengths, to significantly enhance the capabilities of the VideoLLM framework. By strategically using these specialized encoders, we aim to capture a broader spectrum of visual information, enriching VideoLLMs’ understanding of videos.
@@
Combining multiple encoders for multimodal LLMs is gaining attention. Eyes Wide Shut (Tong et al., 2024) explored mixing DINOv2 and CLIP features for LLaVA, but their results signal that mixing features effectively requires investigation. Both Mipha (Zhu et al., 2024) and Prismatic-VLMs (Karamcheti et al., 2024) found that image encoders like CLIP and SigLIP, which are trained using vision-language contrastive loss, surpass other image encoders such as ViT and DINOv2, with SigLIP showing further improvements over CLIP. SPHINX-X (Gao et al., 2024) and SPHINX (Lin et al., 2023b) combines multiple image encoders by concatenating features along the channel dimension, while BRAVE (Kar et al., 2024) concatenates features from multiple encoders sequence-wise, followed by a QFormer with masked modeling. There is also the popular body of research on multimodal LLMs using many modalities including image, video, audio and/or 3D  (Chen et al., 2023; Han et al., 2024; 2023; Li et al., 2023a; Liu et al., 2024a; b; Lyu et al., 2023; Panagopoulou et al., 2023; Su et al., 2023; Sun et al., 2023a; Zhang et al., 2023a). In contrast, this paper dives into the video-language domain, exploring combining multiple image and video encoders and exploiting their structural similarities. Our feature fusion is both performant and efficient in FLOPs, and results in an all-encompassing additive mixture of features which previous works could not achieve.
@@ 3 MERV: Multi-Encoder Representation of Videos
Our goal for MERV is to systematically build a video model that leverages multiple encoders with an LLM to process a video following the LLaVA/PrefixLM (Liu et al., 2023; 2018) paradigm (see Figure 2). Unlike previous works, our focus is not on combining multiple modalities (Bachmann et al., 2022; Zhu et al., 2023), but instead on combining multiple image and video encoders trained on different datasets and objectives. We extensively ablate three key aspects to make this possible: our selection of multiple encoders, i.e., which visual encoders and how many to use (Sec 3.1); how we align the spatio-temporal representations of each encoder to mix the information together, especially in an efficient manner (Sec 3.2); and our implementation efficiencies, from the parallel visual processing to the training recipes (Sec 3.3).
@@ 3.1 Multi-Encoder Feature Extraction
Figure 2: Overview of MERV, a Multi-Encoder Representation of Videos. MERV proceeds in three main stages. First, we feed in our input video into each of visual encoders to get different representations. They are then spatio-temporally aligned before being fused by a cross-attentive mixer. The output is a visual embedding with an additive mix of information from all of the encoders, which is combined with the text query to produce our final generation.
Our final architecture uses four distinct types of models: spatial experts, fine-grained temporal experts, image-language experts, and video-language experts. We found experimentally that our choice of four performed the best across all types of questions, and ablate our choices in Section 4.2. More details about these four encoders and other encoders we considered are in Appendix Table 4.
Spatial expert: DINOv2 (Oquab et al., 2023) is trained using unsupervised learning on local-to-global correspondences in image data. The resulting features have robust object part understanding, as well as semantic and low-level image understanding, but can suffer from poor language grounding.
@@
Video-Language contrastive expert: Finally, our video-language expert is LanguageBind (Zhu et al., 2023). Used by Video-LLaVA (Lin et al., 2023a), LanguageBind is trained through joint multimodal learning between text and multiple modalities, including videos, infrared, and audio, and understands the relationship between video and text and their high-level semantics. We only use the video encoder of LanguageBind.
@@ 3.2 Spatio-Temporally Aligned Representations for Feature Fusion
Our input is a batch of text, image-text, or video-text queries. The visual part of the input, either images $\mathcal{I}$ or videos $\mathcal{V}$, is passed through each of the visual encoders to extract the respective features. Here we describe the detailed care we took in pre-processing to prepare the features for alignment.
@@
Pre-fusion projection. Now we need to achieve spatial alignment among the features. Naïvely combining them is not possible as they all have different spatial shapes, and would also be prohibitively expensive at full resolution. We design a pre-fusion projector to both align and compress them.
@@
Feature fusion strategies. The final part of our pipeline is fusing the multi-encoder information together using cross-attention with learnable queries to additively mix the different representations together. The visual features determine the weights of the linear mixture, which we find sufficient for our task. We use a single randomly initialized query $\mathbf{Q}\in\mathbb{R}^{1\times d}$, keys as $\mathbf{\overline{X}}=[\overline{{\mathbf{x}}}_{1}\;\dots\;\overline{{\mathbf{x}}}_{N}]\in\mathbb{R}^{N\times d_{L}}$, where $\overline{{\mathbf{x}}}_{e}\in\mathbb{R}^{d}$ is each encoder’s features averaged over the sequence dimension $\ell$ for a faster computation, and $N$ the number of encoders, and values as $\mathbf{X}=[{\mathbf{x}}_{1}\;\dots\;{\mathbf{x}}_{N}]\in\mathbb{R}^{N\times\ell\times d}$. We calculate our final unified feature as
@@
The final step is to concatenate the visual embedding and tokenized text together into the LLM. We use the base LLaMA-2 7B model (Touvron et al., 2023b), which we found performs better than the chat model. We test multiple alternate feature fusion strategies and their tradeoffs in Section 4.2.2.
3.3 Implementation Efficiencies
Parallelized visual encoding. At a first glance, using multiple encoders seems to be a large cost to pay when comparing the raw FLOPs and parameters. However, a key benefit of the LLaVA style architecture is that the entire feature extraction and projection pipeline can happen in parallel. To make this possible, we build on top of recent powerful advances in parallel processing for LLMs and use PyTorch’s Fully Sharded Data Parallel (Zhao et al., 2023). As the video encoders themselves are much smaller than the LLM blocks and complete in around the same time, most of the overhead in running four encoders is already covered by having one encoder. We provide some timing numbers in Section 4.2.3 and find that our step time is similar to that of the single-encoder methods.
Our code is built on top of the Prismatic VLM codebase (Karamcheti et al., 2024), which efficiently implements vision-language model (VLM) training. We add the ability to handle videos and an arbitrary number of visual encoders, along with many useful features for training. Our training is efficient for using multiple visual models, completing in under 24 hours using 8 L40-48GB GPUs, and down to 8 hours using 8 H100s. The Video-LLaVA codebase runs Stage 2 in around 38 hours on the same L40 setup and could not easily support multiple encoders in our initial attempts.
MERV frozen and full. Many different recommendations for training LLaVA style models have been made since its inception. This is only made more complicated by the introduction of new datasets with every new VideoLLM architecture, making it difficult to properly determine the best recipe for one’s own setup. We intentionally fix our dataset to be the same as Video-LLaVA’s so we can isolate the impacts of the training setup, from which we find two viable settings. MERV (frozen), which performs only Stage 2 instruction tuning and achieves similar results to the original Video-LLaVA recipe in only 43% of the time, and MERV (full), which unfreezes the LLM during Stage 1 as well for a slight improvement on a few benchmarks. As MERV (frozen) is faster to train with similar performance, we adopt that recipe by default for analysis, and interchangeably use MERV to refer to it for simplicity from here on out. Detailed analysis is provided in Section 4.2.3.
4 Experimental Results
In this section, we show that our method outperforms prior works across standard video benchmarks before moving onto an in-depth analysis of our method and its specializations in the next section.
Datasets and training procedure. Our data mix is the same as Video-LLaVA (Lin et al., 2023a). The Stage 1 data is single-turn concise captioning, with 558k (image, text) pairs from LAION filtered by LLaVA (Liu et al., 2023) and 702k (video, text) pairs from Valley (Luo et al., 2023). The Stage 2 data is multi-turn conversations, detailed captioning and reasoning, with 665k (image, text) pairs from LLaVA and 100k (video, text) instructions from Video-ChatGPT (Maaz et al., 2023).
4.1 Comparison to State of the Art
@@
LLaMA-Adapter (Zhang et al., 2023b) 54.9 3.1 43.8 2.7 - - - 34.2 2.7 - - -
Video-LLaMA (Zhang et al., 2023a) 51.6 2.5 29.6 1.8 - - - 12.4 1.1 - - -
Video-ChatGPT (Maaz et al., 2023) 64.9 3.3 49.3 2.8 - - - 35.2 2.7 - - -
SeViLA (Yu et al., 2024) - - - - - - 46.2 - - 63.6 64.4 38.2
LLaMA-VID-7B* (Li et al., 2023e) 69.30 3.74 57.84 3.24 51.31 3.26 41.64 46.45 3.22 60.61 57.65 37.43
LLaMA-VID-13B* (Li et al., 2023e) 70.25 3.77 58.58 3.26 51.26 3.26 41.54 46.79 3.23 60.03 61.98 41.33
@@
Video-LLaVA* (Lin et al., 2023a) 67.74 3.69 56.90 3.18 47.99 3.17 44.22 47.08 3.27 59.61 61.21 37.66
MERV (frozen) 70.97 3.76 59.03 3.25 51.1 3.26 46.21 50.87 3.34 63.09 58.66 42.28

### v2

@@ Abstract
Recent advances in vision backbones have yielded powerful and diverse visual and video encoders. Yet, current Video Large Language Models encode visual inputs using an encoder from a single backbone family, limiting the amount and type of visual information they can process. We propose MERV, a Multi-Encoder Video Representation, which utilizes multiple encoders for a comprehensive video representation. To optimize heterogeneous features from a broad spectrum of encoders and ensure efficient and coherent feature integration, MERV first aligns encoder features spatio-temporally, then projects them into a unified structure, and finally fuses them through cross-attention. Under fair comparison, MERV achieves up to 4.62% higher accuracy than its base model, while introducing minimal extra parameters and training faster than equivalent single-encoder methods after parallelizing visual processing. Qualitative analysis shows MERV successfully captures and integrates domain knowledge from each encoder, opening new possibilities for scaling enhanced video understanding.
Machine Learning, Video Understanding, LLMs, ICML
@@ 1 Introduction
Inspired by the sophisticated reasoning abilities of recent Large Language Models (LLMs) (Chiang et al., 2023; Chowdhery et al., 2023; OpenAI, 2023), researchers have focused on using them in many other domains to great success. The video counterparts, known as Video Large Language Models (VideoLLMs) (Bain et al., 2021; Li et al., 2023c; Lin et al., 2024; Luo et al., 2023; Maaz et al., 2024; Yu et al., 2024b), connect pre-trained vision encoders to LLMs by training a modality bridge from the vision space to the language space, allowing for reasoning to happen in the highly expressive language domain.
@@
Figure 1: Examples where a single encoder model is the only model to correctly answer Perception Test questions (Pătrăucean et al., 2023), while MERV can correctly answer all types.
Most multimodal LLMs, such as LLaVA (Liu et al., 2023) for images and Video-LLaVA (Lin et al., 2024) for videos, opt for contrastively pre-trained encoders like CLIP (Radford et al., 2021) and LanguageBind (Zhu et al., 2024a). Their vision-language pre-training naturally lends itself as a bridge between the vision input and the LLM, circumventing the need to train heavy vision-language alignment modules like a QFormer (Li et al., 2022). These encoders are almost always pre-trained separately and vary in architecture, training data, and optimization strategy. Consequently, the features extracted by these encoders exhibit unique characteristics, each with inherent strengths and limitations. Contrastive encoders like CLIP may be better suited with their multimodal semantic alignment, but are inferior to models such as DINOv2 (Oquab et al., 2023) at fine-grained object level understanding. They also fail to take advantage of models trained specifically on videos, such as ViViT (Arnab et al., 2021). Despite this clear tension between vision backbones, previous research in VideoLLMs has relied on only one vision encoder for visual processing as one was thought to be sufficient for visual understanding, and already difficult enough to achieve vision-language alignment with. Any more encoders was unnecessary and not an effective tradeoff of runtime for compute.
Figure 2: MERV architecture and performance. (Left) MERV proceeds in three main stages. First, we feed in our input video into each of visual encoders to get different representations. They are then spatio-temporally aligned before being fused by a cross-attentive mixer. The output is a visual embedding with an additive mix of information from all the encoders, which is combined with the text query to generate the result. (Right) We only compare MERV to the 7B prior models trained with comparable training data mixes. MERV is on-par or better than these single-encoder prior works, while deliver considerable gains over Video-LLaVA (Lin et al., 2024), from which MERV was initially adapted to form a Multi-Encoder Video Representation.
@@
In this paper, we argue that this choice to not use multiple encoders in existing VideoLLMs unnecessarily restricts their capabilities. For example, in Figure 1 we can see cases where only one of four different single-encoder models answers a given question correctly. While simple scene descriptions can be answered by image-level models, other questions require temporal and action-level comprehension, benefiting from features encoded with video models like ViViT (Arnab et al., 2021). Consequently, the reasoning capabilities of these VideoLLMs are directly limited by the inherent weaknesses of their respective pre-trained encoders. Therefore, employing multiple encoders could allow us to complement one encoder’s weaknesses with another encoder’s strengths. It also offers a cheap way to broaden the input distribution, leveraging different encoder’s training mixtures. The wide adoption of the LLaVA paradigm is also indication that vision-language alignment is simple to achieve, even without language-aware vision models.
@@
We propose MERV, a Multi-Encoder Representation of Videos, as a new method for integrating multiple visual encoders into a single VideoLLM using a cross-attentive encoder mixer for fusing representations. We introduce a spatio-temporally aligned representation for mixing the information from multiple types of visual encoders. Given the computational complexity of video tasks, we carefully experiment with optimization strategies and parallelizing the visual experts, allowing us to combine four distinct visual encoders with minimal computational overhead. Our frozen method outperforms all individual encoder methods, is up to 4.6% better than prior works (Lin et al., 2024) on video reasoning benchmarks, i.e., from 37.66% to 42.28% on TVQA (Lei et al., 2018), and on par with SeViLA (Yu et al., 2024b) on Perception Test (Pătrăucean et al., 2023), a challenging perception and reasoning diagnostic for video models. Finetuning the full model improves MERV past SeViLA (Yu et al., 2024b) by 2.2%, from 46.2% to 48.4%. Finally, we do a detailed qualitative study of our model’s capabilities on the Something-Something v2 dataset (Goyal et al., 2017). We show that MERV can accurately capture both contrastive encoders’ (Zhai et al., 2023; Zhu et al., 2024a) strengths on general vision-language understanding, as well as ViViT’s (Arnab et al., 2021) specialty on temporally-sensitive tasks (e.g. distinguishing pushing left vs. right), without trading off performance between these specializations as single encoder models do.  [^1] Our code and pretrained weights are available at https://github.com/princetonvisualai/merv.
@@ 2 Related Works
VideoLLMs build upon the powerful reasoning capabilities of LLMs by utilizing them as language decoders to enable instruction-followed video understanding. Key advancements include VideoChat (Li et al., 2023c) and Video-LLaMA (Zhang et al., 2023) for chat-based video understanding, LLaMA-Adapter (Zhang et al., 2024b) for pre-alignment, Valley (Luo et al., 2023) with multilingual LLMs, InternVideo (Wang et al., 2022) with a dedicated video encoder training phase, and Video-ChatGPT (Maaz et al., 2024) combining video-adapted encoders with LLMs. GPT4Video (Wang et al., 2023) supports video understanding and generation, while MovieChat (Song et al., 2024) focuses on long video comprehension. Models like Chat-UniVi (Jin et al., 2024), LLaMA-VID (Li et al., 2024b), and VTM  (Lee et al., 2024) optimize token usage for video representation. Other notable models include Vamos (Wang et al., 2024b), which flexibly uses visual embeddings, action labels, and video captions as input; VideoChat2 (Li et al., 2024a), developed through three-stage progressive training; Video-LLaVA (Lin et al., 2024), which aligns image and video representations before projecting them to the LLM space; and VideoPrism (Zhao et al., 2024), which also further trains a video encoder through masked distillation. Specialized models like VTimeLLM (Huang et al., 2024) focus on fine-grained video moment understanding and time-bound reasoning, while models like Elysium (Wang et al., 2024a) and Merlin (Yu et al., 2024a) can predict object trajectories. SeViLA (Yu et al., 2024b) uses LLM for frame localizer of the video for multiple-choice tasks. Finally, recently LLaVA-Hound-DPO (Zhang et al., 2024a) explored using DPO and a higher quality training set for better instruction following. Distinct from these aforementioned works, our approach centers on utilizing a diverse array of visual and video encoders, which are RGB-based but coming from different visual backbone families, each with its own unique strengths, to significantly enhance the capabilities of the VideoLLM framework. By strategically using these specialized encoders, we aim to capture a broader spectrum of visual information, enriching VideoLLMs’ understanding of videos.
@@
Combining multiple encoders for multimodal LLMs is gaining attention. Eyes Wide Shut (Tong et al., 2024b) explored mixing DINOv2 and CLIP features for LLaVA, but their results signal that mixing features effectively requires investigation. Both Mipha (Zhu et al., 2024b) and Prismatic-VLMs (Karamcheti et al., 2024) found that image encoders like CLIP and SigLIP, which are trained using vision-language contrastive loss, surpass other image encoders such as ViT and DINOv2, with SigLIP showing further improvements over CLIP. SPHINX-X (Gao et al., 2024) and SPHINX (Lin et al., 2023) combines multiple image encoders by concatenating features along the channel dimension, while BRAVE (Kar et al., 2024) concatenates features from multiple encoders sequence-wise, followed by a QFormer with masked modeling. More recently, Cambrian-1 (Tong et al., 2024a) creates a multimodal LLM which spatially aligns their inputs across different resolutions. This also reproduces our finding of the importance of spatio-temporal alignment. There is also the popular body of research on multimodal LLMs using many modalities including image, video, audio and/or 3D  (Chen et al., 2023; Han et al., 2023; Li et al., 2023a; Lyu et al., 2023; Panagopoulou et al., 2023; Su et al., 2023; Sun et al., 2023a; Zhang et al., 2023; Liu et al., 2024a; Han et al., 2024; Liu et al., 2024b; Jain et al., 2024). In contrast, this paper dives into the video-language domain, exploring combining multiple image and video encoders and exploiting their structural similarities. Our multi-encoder feature alignment and fusion are both performant and efficient in FLOPs, and results in an all-encompassing additive mixture of features which previous works could not achieve.
@@ 3 MERV: Multi-Encoder Representation of Videos
Our goal for MERV is to systematically build a video model that leverages multiple encoders with an LLM to process a video following the LLaVA/PrefixLM (Liu et al., 2023, 2018) paradigm (see Figure 2). Unlike previous works, our focus is not on combining multiple encoders of different modalities that make use of additional information (depth, audio, etc) (Bachmann et al., 2022; Zhu et al., 2024a), but instead focus on RGB-encoders trained on different datasets and objectives that offer different visual understanding. We extensively ablate three key aspects to make this possible: our selection of multiple encoders, i.e., which visual encoders and how many to use (Sec 3.1); how we align the spatio-temporal representations of each encoder to mix the information together, especially in an efficient manner (Sec 3.2); and our implementation efficiencies, from the parallel visual processing to the training recipes (Sec 3.3).
@@ 3.1 Multi-Encoder Feature Extraction
Our final architecture uses four distinct types of models: spatial experts, fine-grained temporal experts, image-language experts, and video-language experts. We found experimentally that our choice of four performed the best across all types of questions, and ablate our choices in Sec. 4.2. More details about these four encoders are in Appendix Table 4.
Spatial expert: DINOv2 (Oquab et al., 2023) is trained using unsupervised learning on local-to-global correspondences in image data. The resulting features have robust object part understanding, as well as semantic image understanding, but can suffer from poor language grounding.
@@
Video-Language contrastive expert: Finally, our video-language expert is LanguageBind (Zhu et al., 2024a). Used by Video-LLaVA (Lin et al., 2024), LanguageBind is trained through joint multimodal learning between text and multiple modalities (videos, infrared, etc.,) and understands the relationship between video and text and their high-level semantics. We only use the video encoder of LanguageBind.
@@ 3.2 Spatio-Temporally Aligned Representations for Feature Fusion
Our input is a batch of text, image-text, or video-text queries. The visual part of the input, either images or videos, is passed through each of the visual encoders to extract the respective features. Here we describe the detailed care we took in pre-processing to prepare the features for alignment.
@@
Pre-fusion projection. Now we need to achieve spatial alignment among the features. Naïvely combining them will not work as they all have different spatial shapes. Using the full resolution features would also be prohibitively expensive. We design a pre-fusion projector to tackle both issues by aligning and compressing the features.
@@
Feature fusion strategies. The final part of our pipeline is fusing the multi-encoder information together using cross-attention with learnable queries to additively mix the different representations together. The visual features determine the weights of the linear mixture, which we find sufficient for our task. We use a single randomly initialized query $\mathbf{Q}\in\mathbb{R}^{1\times d}$, keys as $\mathbf{\overline{X}}=[\overline{{\mathbf{x}}}_{1}\;\dots\;\overline{{\mathbf{x}}}_{N}]\in\mathbb{R}^{N\times d}$, where $\overline{{\mathbf{x}}}_{e}\in\mathbb{R}^{d}$ is each encoder’s features (after Eq. 1) averaged over the sequence dimension $\ell$ for a faster computation, and $N$ the number of encoders, and values as $\mathbf{X}=[{\mathbf{x}}_{1}\;\dots\;{\mathbf{x}}_{N}]\in\mathbb{R}^{N\times\ell\times d}$. We calculate our final unified feature as
@@
The final step is to concatenate the visual embedding and tokenized text together into the LLM. We use the base LLaMA-2 7B model (Touvron et al., 2023b), which we found performs better than the chat model. We test multiple alternate feature fusion strategies in Section 4.2.2.
Table 1: Comparison of different multimodal LLMs on video reasoning benchmarks. We employ ChatGPT to evaluate performance following Video-ChatGPT where applicable (version gpt-3.5-turbo-0613). * denotes our evaluation of using the author provided checkpoint. The first five datasets were used as evaluation test sets during development; the last three were held-out for our final evaluation.
@@
LLaMA-Adapter (Zhang et al., 2024b) 54.9 3.1 43.8 2.7 - - - 34.2 2.7 - - -
Video-LLaMA (Zhang et al., 2023) 51.6 2.5 29.6 1.8 - - - 12.4 1.1 - - -
Video-ChatGPT (Maaz et al., 2024) 64.9 3.3 49.3 2.8 - - - 35.2 2.7 - - -
SeViLA (Yu et al., 2024b) - - - - - - 46.2 - - 63.6 64.4 38.2
LLaMA-VID-7B* (Li et al., 2024b) 69.30 3.74 57.84 3.24 51.31 3.26 41.64 46.45 3.22 60.61 57.65 37.43
LLaMA-VID-13B* (Li et al., 2024b) 70.25 3.77 58.58 3.26 51.26 3.26 41.54 46.79 3.23 60.03 61.98 41.33
@@
Video-LLaVA* (Lin et al., 2024) 67.74 3.69 56.90 3.18 47.99 3.17 44.22 47.08 3.27 59.61 61.21 37.66
MERV (frozen) / MERV 70.97 3.76 59.03 3.25 51.1 3.26 46.21 50.87 3.34 63.09 58.66 42.28